urrently, the stock price is $50. Over each of the next two 1-yr periods it is expected to go up by 20% or down by 20%. The risk-free rate is 5% per annum with continuous compounding.

What is the value of a 2-yr European call option with a strike price of $61? Round to the nearest cent. For example, if your answer is $12.345, then enter 12.35. Margin of error: +/- 0.10.
Answer #1

Step 5 Binomial Tree Pricing Step 1 Step 2 Scenario price-Spot price*(U)^2=50*1.2^2=72 Payoff HH=Max(Scenario price-Strike pr

therefore value of call option = 3.93